Amedical researcher says that less than 26% of adits in aontain there enough evidence to support the researchers claim? Complete parts (a) through (e) below. (a) lidently the claim and state Ho and H What is the claim? OA. Less than26% of all adults are smokers o untry are sman. anno sample o 250 adde fom that ountry. 168% say t they n Na- s. e Less than 20% of adults in the country aro smokers. Exactly 16.8% of al adults are smokers. Exactly 16.8% of aduts in the country are smokers B. OC. D. identity Ho and H (Type integers or decimals)
media%2F333%2F333cf91d-23ac-4f2e-bc27-22
Answer #1

H0: P= 0.2

HA: P<0.2

Since it is left tailed test critical value of Z= -1.645

Rejection region for left tailed test is

Z<-1.645

Standardized Z statistic is

Z=p-P/sqrt(PQ/n)

Z= 0.168-0.2/sqrt(0.2*0.8/250)

Z= -0.032/sqrt(0.00064)

Z= -0.032/0.0253

Z=-1.26

Since Calculated value of Z is greater than critical value of Z for left tailed test, we fail to reject NULL HYPOTHESIS H0 at 0.05 level of significance.

Fail to reject H0. There is no enough evidence to support the reserachres's claim.